MODS2 Commissioning: Night 2

Telescope Manager: Michael Wagner
Telescope Operator: David Gonzalez Huerta
Instrument Specialist: Jenny Power
Observers: Mark Wagner, Olga Kuhn (in Tucson)

Sunday it was found that one hard disk in the MODS2 Red IC was not working. Replacement was shipped, brought to mountain today and installed. M2.RC is now working.

Monday was to be night 1 of MODS2 Commissioning - forfeited bc of disk problems. This is technically night 2.

execMODS2 ~modseng/modsScripts/Support/modsWake.pro

execMODS2 ~modseng/modsScripts/Support/sieveSnap.pro
  • mods2(r/b).20141230.0001.fits Blue image has jump. Repeat
  • mods2(r/b).20141230.0002.fits Both Blue and Red images OK now.

Switch root name to 20141231
execMODS2 ~modseng/modsScripts/modsCalib/bias3K.cal

Too windy to open. Running Closed Dome tests:

00:58 CCDROTime
  • b 6-45
  • r 6-45
  • completed around 01:37

01:40 execMODS2 ~modseng/Commissioning/ClosedDome/photTran_b.cal
  • b 46... 75
    • mods2b.20141231.0050.fits shows horizontal banding - sequencer or Blue HEB problem? But it is intermittent (see 52, which looks normal). Bad region on earlier images different from quadrant to quadrant, but 53 is again bad, and on this one, same rows affected for all quadrants. 56 is also bad, again same rows for all 4 quadrants --- this suggests a sequencer problem.
  • in all, about 7 of the 28, or 1/4, showed the sequencer problems. Skipping to the spectral mapping scripts for now.

02:55 execMODS2 ~modseng/Commissioning/ClosedDome/pinholeSlit.cal
  • b 76 - 81 --- all appear free of sequencer problems
  • r 46 - 52

03:13 execMODS2 ~modseng/Commissioning/ClosedDome/sudokuMask.cal
  • b 82 - 87
  • r 53 - 59
    • r 53 --- look at the horizontal streaking from the pinholes at the left, this is like the sieveSnaps at the beginning of the night.
    • b 83 (QTHLamp1), 87 (ArLamp) shows sequencer problem, will repeat sequence.
  • b 88 - 93
    • b 89 (QTHLamp1) and 90 (QTHLamp2) are bad because of sequencer problem.
    • b 92 (KrXeLamp) shows image jumps (an imcs problem? see 1st blue sieveSnap this afternoon, 20141230.0001) & r 60-66

03:55 execMODS2 ~modseng/Commissioning/ClosedDome/pr_sudoku.cal
  • Script stopped bc blue imcs failed to lock, it locked after the =R=etry.
  • b 94 - 97
    • b94 (Ne+Hg) and b96 (Kr)
  • r 67 - 70

04:16 execMODS2 ~modseng/Commissioning/ClosedDome/wiggle.cal
  • b 98 - 101
    • all four blue images are bad.
  • r 71 - 74

04:32 execMODS2 ~modseng/Commissioning/ClosedDome/gg495.cal
  • r 75-76

04:38 execMODS2 ~modseng/Commissioning/ClosedDome/photTran_r.cal
  • r 77- 100

05:24 execMODS2 ~modseng/Commissioning/ClosedDome/darkTest.cal
  • b 102 - 119
    • a few show amplifier glow in each of the 4 corners, e.g. 102, 115
    • a hot pixel at (1991,386) where Y=386 is the quadrant boundary for the 4x4 binning used. This hot pixel affects both the column below and row to the left. See 102 for example.
  • r 101 - 118
    • r 112-113 show amplifier glow
    • r 111 also shows some dark rows on bottom right quadrant; same effect but on upper left quadrant for r 112. Seen earlier on a Sudoku or Pinhole spectrum.
  • amplifier glow very strong on both blue and red channel 20-min exposures, see r 116 and b 117

07:30 Going to cycle power to the HEB and the IC program to see whether this will have any effect on the blue images.
  • Log in to KVM, connect to M2.BC, quit from IC
  • isisCmd --mods2 m2.ie util heb_b off
  • isisCmd --mods2 m2.ie util heb_b on
  • restart IC on M2.BC

07:35 Repeating wiggle.cal, where all 4 blue images were bad earlier (mods2b.20141231.0098 - 0101)
  • b 120-123 all bad again.
    • QTH1 ND1.5 5sec
    • QTH1 UG5 4sec
  • r 119-122
  • try one more b with ND1.5 and 2-sec exposure - also bad.

08:00 We have not done the leakTest nor the linTest scripts for b and r. But a significant fraction of the blue images are bad due to what looks like a problem with the sequencer (but should be confirmed). We are calling it a night.

-- %USERSIG{OlgaKuhn - 2014-12-30}%

Comments

 
Topic revision: r3 - 31 Dec 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback